OpenStudy (anonymous):

Two concentric spheres have radii of 5" and 6". Find the volume of the space between them.

OpenStudy (anonymous):

To find the volume of the space between two concentric spheres, you would calculate the volume of both spheres and then subtract the volume of the smaller sphere from the volume of the large sphere. Hint: Volume of a sphere V=43πr3

OpenStudy (mrnood):

It doesn't matter if they are concentric - provided the smaller is totally enclosed inside the larger then the volume between them is simply the difference between their 2 volumes. Volume of a sphere =\[\frac{ 4 }{ 3 } \pi r ^{3}\]

OpenStudy (anonymous):

MrNood we gave pretty much same answer.

OpenStudy (mrnood):

@mjoxprm360 Yes sorry - I was typing whilst yours appeared

OpenStudy (anonymous):

Volume of a Sphere = 4/3 π r³ Volume of the space in between = > 4/3 π (6)³ - 4/3 π (5)³ = 121 π in³

OpenStudy (anonymous):

@MrNood I wasn't complaining I am happy you got it right as well.

OpenStudy (anonymous):

i.e 46.01 cubic units

OpenStudy (mrnood):

However - I would discourage you from completing the answer! It is not too much ot expect the OP to do the work once the process has been explained!

OpenStudy (anonymous):

Yes I know :)

OpenStudy (mrnood):

@chakradhar The units are given as inches - however - your answer is incorrect

OpenStudy (mrnood):

@chakradhar I see oyu have calculated using r^2 instead of r^3

OpenStudy (anonymous):

Absolutely

OpenStudy (anonymous):

yes...sryy dudes